# Realistic models of general-relativistic differentially rotating stars

### v1

@@ 5.3 Comparison of various differential-rotation laws
Figure 13: The same as in Fig. 2, but for all of the differential-rotation laws considered here: $j$-constant law (violet line), U9 law (green and light-blue lines) the U9+ law (orange line) and the new CR law (red line); see also Figs. 5, 10, and 11. The Keplerian limit of differentially rotating models obeying the $j$-constant law for $\hat{A}=3.33$, i.e., $\Omega=\Omega_{\rm K,dr}$, is shown with a pink dashed line; the solid filled circles mark the stellar models shown in Fig. 15.
@@ 6 Conclusions
In particular, we have analysed the results of three, state-of-the-art BNS merger simulations performed with different and realistic EOSs and mass ratios, and have extracted from them the quasi-stationary rotational properties of the remnants, both in terms of azimuthally- and time-averaged profiles of the angular velocity and of the specific angular momentum. This analysis has been useful to study the existence of universal features in the angular-velocity profile and to point out a simple criterion to recognise the transition between the “core” of the remnant and the “disc”, that is, the portion of the HMNS whose angular-velocity profile follows a Keplerian behaviour. More specifically, it was possible to realise that in the remnant core the angular velocity exhibits a robust and generic linear behaviour in terms of the logarithm of the rest-mass density. Furthermore, the departure from this linear behaviour also corresponds to the location in the remnant where the Keplerian flow starts, so that it is simple to mark the beginning of the remnant disc as the location where the angular velocity is no longer linearly proportional to the logarithm of the rest-mass density. Interestingly, this position does not correspond to the maximum of the angular velocity, but is located at larger radii.
@@
The systematic and comprehensive analysis presented here can be extended and improved in a number of ways. First, by exploring the post-merger properties of remnants produced with other EOSs so as to validate the robustness of the proposed U9+ and CR laws. Particularly important in this respect will be exploration of a broader range of mass ratios as they seem to exhibit somewhat different differential-rotation laws. Second, by improving the convergence properties of Hydro-RNS and hence compute the Keplerian line of differentially rotating models that so far in the literature has been computed only for the simpler $j$-constant differential-rotation law. Third, compute the location of the neutral-stability line (Takami et al., 2011) and determine its location with respect to the turning-point stability line. Finally, use numerical simulations of dynamical spacetimes to explore the equilibrium properties of the turning-point and neutral stability lines to determine necessary and sufficient conditions for collapse to rotating black holes, thus confirming also for more realistic differentially rotating stars the analysis carried out for the $j$-constant law (Weih et al., 2018; Szewczyk et al., 2023). We plan to explore some of these topics in future works.
@@ Appendix A On the U8 models and the DD2 simulation
Appendix B Computing equilibria with the CR law
This Appendix is dedicated to the numerical strategy we have implemented in order to obtain numerical solutions with the new and more complex CR differential rotation law. We recall that it has free parameters $\beta$, $\gamma,A,\lambda_{1}$, and $j_{\rm max}/j_{e}$ which, together, control the shape of the profile. In particular, the parameter $\lambda_{1}$ regulates the maximum angular velocity and can be read of from the ratios of the angular velocities from the merger-profiles. The parameter $\beta$ controls instead the slope of the angular velocity for angular momenta smaller than $j_{\rm max}$, which tends to become constant (linear behaviour) for smaller values of $\beta$. The parameters $A$ and $\gamma$ are responsible for the part with angular momenta larger than $j_{\rm max}$, with $\gamma$ controlling how far from the maximum the decreasing part is extended, and $A$ regulating the slope of the angular velocity in the Keplerian region. Note that larger values of the ratio $\gamma/A$ yield profiles that extend to larger angular momenta.
While the CR law provides a more accurate representation of the remnant rotational properties, it also offers additional challenges during the numerical calculation and can easily lead to a loss of convergence in the iterative procedure employed by Hydro-RNS in computing a solution. In particular, for large values of $\lambda_{1}=\Omega_{\rm max}/\Omega_{c}$, the range of possible guesses for $r_{e}$ and of the corresponding solutions of $\Omega_{e}$ grows substantially. In addition, if the guess for $r_{e}$ is rather different from the one for which a solution exists, the code fails to converge because no solution with the guessed $\Omega_{e}$ can be found with the Euler equation. It is in principle possible to manually control the iterations of $r_{e}$ and to extend the range of search of $\Omega_{e}$ so that a solution is found, but this is tedious and not robust.
However, an approach that tackles these problems consists “blending” the solution with the new CR rotation law with the solution obtained with the U9 law, which is less accurate but less challenging in the regions of large specific angular momentum. Hence, we introduce a smooth transition from the converging U9 rotation law to the new CR law via a simple linear combination
$\Omega(j)=\alpha\ \Omega_{\rm U9+}(j)+\beta\ \Omega_{\rm CR}(j)\,,$ (63)
where $\alpha,\beta\in[0,1]$ and $\alpha+\beta=1$. Hence, starting with $\alpha=1,\,\beta=0$ and progressively marching to a situation in which $\alpha=0,\,\beta=1$, the Hydro-RNS code can find equilibrium solutions with axis ratios that are larger than those allowed by the U9+ rotation law. Clearly, the strategy adopted in the blending (63) with the U9 law can be employed also with other differential-rotation laws.
Figure 17 illustrates the result of this approach by showing how an equilibrium solution with the CR law with $j_{\rm max}/j_{e}=0.82,\beta=0.95,\gamma=1.8,A=20,\lambda_{1}=3$ (red thick line) can be obtained from an initial U9+ law with $p=1,\xi=1,\lambda_{1}=2,\lambda_{2}=1.6$ (blue thick line) via a sequence of intermediate solutions which represent the various flavours of the blending  (63) (blue and red thin lines). The final solution is compared with the angular velocity profile of the DD2 BNS simulation (black thick line) at $t-t_{\rm mer}=20\,{\rm ms}$, showing the superior accuracy of the new CR law.

### v2

@@ 5.3 Comparison of various differential-rotation laws
Figure 14: The same as in Fig. 2, but for all of the differential-rotation laws considered here: $j$-constant law (violet line), U9 law (green and light-blue lines) the U9+ law (orange line) and the new CR law (red line); see also Figs. 5, 10, and 12. The Keplerian limit of differentially rotating models obeying the $j$-constant law for $\hat{A}=3.33$, i.e., $\Omega=\Omega_{\rm K,dr}$, is shown with a pink dashed line; the solid filled circles mark the stellar models shown in Fig. 16.
@@ 6 Conclusions
In particular, we have analysed the results of three, state-of-the-art BNS merger simulations performed with different and realistic EOSs and mass ratios, and have extracted from them the quasi-stationary rotational properties of the remnants, both in terms of azimuthally- and time-averaged profiles of the angular velocity and of the specific angular momentum. This analysis has been useful to study the existence of universal features in the angular-velocity profile and to point out a simple criterion to recognise the transition between the “core” of the remnant and the “disc”, that is, the portion of the HMNS whose angular-velocity profile follows a Keplerian behaviour. More specifically, it was possible to realise that in the remnant disc the angular velocity exhibits a robust and generic linear behaviour in terms of the logarithm of the rest-mass density. Furthermore, the departure from this linear behaviour also corresponds to the location in the remnant where the Keplerian flow starts, so that it is simple to mark the beginning of the remnant disc as the location where the angular velocity is no longer linearly proportional to the logarithm of the rest-mass density. Interestingly, this position does not correspond to the maximum of the angular velocity, but is located at larger radii.
@@
The systematic and comprehensive analysis presented here can be extended and improved in a number of ways. First, by exploring the post-merger properties of remnants produced with other EOSs so as to validate the robustness of the proposed U9+ and CR laws. Particularly important in this respect will be exploration of a broader range of mass ratios as they seem to exhibit somewhat different differential-rotation laws. Second, by improving the convergence properties of Hydro-RNS and hence compute the Keplerian line of differentially rotating models that so far in the literature has been computed only for the simpler $j$-constant differential-rotation law. Third, compute the location of the neutral-stability line (Takami et al., 2011) and determine its location with respect to the turning-point stability line. Finally, use numerical simulations of dynamical spacetimes to explore the equilibrium properties of the turning-point and neutral-stability lines, to determine necessary and sufficient conditions for collapse to rotating black holes, thus confirming also for more realistic differentially rotating stars the analysis carried out for the $j$-constant law (Weih et al., 2018; Szewczyk et al., 2023). We plan to explore some of these topics in future works.